### v2

# Imperial College London, London, UK

Abstract
We apply logic-based machine learning techniques to facilitate cellular engineering and drive biological discovery, based on comprehensive databases of metabolic processes called genome-scale metabolic network models (GEMs). Predicted host behaviours are not always correctly described by GEMs. Learning the intricate genetic interactions within GEMs presents computational and empirical challenges. To address these, we describe a novel approach called Boolean Matrix Logic Programming (BMLP) by leveraging boolean matrices to evaluate large logic programs. We introduce a new system, $BMLP_{active}$, which efficiently explores the genomic hypothesis space by guiding informative experimentation through active learning. In contrast to sub-symbolic methods, $BMLP_{active}$ encodes a state-of-the-art GEM of a widely accepted bacterial host in an interpretable and logical representation using datalog logic programs. Notably, $BMLP_{active}$ can successfully learn the interaction between a gene pair with fewer training examples than random experimentation, overcoming the increase in experimental design space. $BMLP_{active}$ enables rapid optimisation of metabolic models and offers a realistic approach to a self-driving lab for microbial engineering.
Keywords:
Logic Programming Boolean Matrices Systems Biology Genome-Scale Metabolic Networks.
1 Introduction
Our study explores the applicability of abductive reasoning and active learning to extensive biological systems. We advance the scope of logic-based gene function learning by looking at Escherichia coli (E. coli). The genome-scale metabolic model (GEM), iML1515 [19], encompasses 1515 genes and 2719 metabolic reactions of the E. coli strain K-12 MG1655, a versatile host organism for metabolic engineering to produce specific compounds. In iML1515, interactions between gene pairs have been identified as a key error source [3]. Learning these interactions requires exploring a combinatorial hypothesis space. Given that biological relationships are commonly described logically, Inductive Logic Programming (ILP) [20] is particularly adept at operating on biological knowledge bases. We propose Boolean Matrix Logic Programming (BMLP) where we employ boolean matrices as underlying bottom-up evaluation mechanisms of dyadic datalog programs to scale to state-of-the-art GEM iML1515. To learn from a combinatorial hypothesis space with high data efficiency, we create an active learning system $BMLP_{active}$ using a BMLP module called BMLP-IE to learn digenic interaction, requiring 90% fewer experiments than random experimentation. The integration of abductive reasoning and active learning via ILP was successfully demonstrated in the context of biological discovery by the prominent Robot Scientist [14]. The Robot Scientist performed active learning by strategically selecting key experiments to achieve more data- and cost-effective gene function learning than random experimentation. However, this demonstration was limited to only 17 genes in the aromatic amino acid pathway of yeast. Notably, the iML1515 model represents a 90-fold increase in complexity compared to the aromatic amino acid pathway investigated by the Robot Scientist. Our system $BMLP_{active}$ is the first logic programming system to ever learn digenic interaction on the genome scale from GEMs.
Figure 1: Certain genetic mutations would block pathways, causing cells to die (positive label). $BMLP_{active}$ finds a gene-reaction association hypothesis to explain the pathway blockage and lethality. It encodes the GEM iML1515 as boolean matrices and uses BMLP-IE to classify the genetic mutation experiment labels for every hypothesis. It consults a data source to request ground truth labels. $BMLP_{active}$ iteratively refutes hypotheses inconsistent with ground truth labels.
2 Related work
Computational systems [16, 30, 4, 12, 24] were primarily designed to formulate symbolic hypotheses from experimental results. One cannot directly employ these systems for experimental planning. In contrast, the Robot Scientist [14, 15] automatically proposed explanatory hypotheses, actively devised validation experiments and performed experiments using laboratory robotics. Experiments were actively selected to minimise the expected cost of experimentation for learning gene functions in yeast [5]. In comparison, we apply $BMLP_{active}$ to a GEM, which is significantly larger than the models in the Robot Scientist [14, 15].
Our BMLP approach uses boolean matrices for iterative bottom-up computation. Obtaining the least Herbrand model of linear recursive datalog programs can be reduced to computing the transitive closure of boolean matrices [23, 8]. Fischer and Meyer [10] and Ioannidis [13] showed a divide-and-conquer boolean matrix computation technique by viewing relational databases as graphs. An ILP system called DeepLog [21] employed boolean matrices for choosing optimal background knowledge to derive a compact bottom clause. A great body of work [17, 11, 27, 6, 29, 28] only studied the approximation of logic program evaluation in tensor spaces. Solving certain recursive programs is difficult whereas these can be solved by iterative bottom-up evaluations [27].
3 Modelling a GEM with a Petri net
In an auxotrophic mutant experiment, a gene is removed from a cell. Some key metabolic reactions might no longer be viable due to the removal, leading to insufficient cell growth. We model such metabolic mechanisms via GEMs. A GEM contains biochemical reactions of substances, which is often represented by Petri nets [26]. A GEM is a Petri net in the sense that reactions are transitions (edges in a Petri net), and reactants and products are places (nodes in a Petri net). A reaction involves chemical substances $x_{i}$ and $y_{j}$:
[…]
We associate binary phenotypic effects, e.g. reduced cell growth between wild-type and auxotrophic mutants, with reachability to key substances in the metabolic network. For this goal, we look at Petri nets with restrictions, namely one-bounded elementary nets (OEN) where place nodes are marked with at most one token [25] (Example 1). The reachability problem is evaluated as a recursive datalog program with arity two. Each argument of our transformed datalog program is a concatenation of multiple constants, e.g. $h2\_o2$. This transformation with constant concatenation generalises the structure of a Petri net into a hypergraph [2], a directed graph where an edge can connect to any number of nodes. To construct this datalog program, we consider a hypergraph with the same bipartite incidence structure between substances and reactions.
Example 1
$\mathcal{P}_{1}$ represents a metabolic network as an OEN and hypergraph. Transitions $t_{1}$, $t_{2}$ and $t_{3}$ are described by $reaction\_1$, $reaction\_2$ and $reaction\_3$ clauses. $c1\_c2$ and $c3\_c4$ are two hypernodes. Both graphs show that the node $c_{5}$ is reachable from $\{c_{1},c_{2}\}$ via $t_{1}$ and $t_{2}$. We can compute these by evaluating the groundings of $pathway(m1,Y)$ in the Herbrand model which gives us $pathway(m1,c2)$, $pathway(m1,c3\_c4)$ and $pathway(m1,c5)$.
$\mathcal{P}_{1}:\left\{\begin{aligned} metabolites(m1,c1\_c2).&\\ metabolites(m1,c1).&\\ metabolites(m1,c2).&\\ reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ reaction(X,Y)\leftarrow metabolites(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
[Uncaptioned image][Uncaptioned image]
4 Boolean matrix logic programming
We propose the Boolean Matrix Logic Programming (BMLP) problem. In contrast to traditional logic program evaluation, BMLP uses boolean matrices to evaluate recursive datalog programs with arity at most two and at most two body literals, namely the $H_{2}^{2}$ program class [22].
Definition 1 (Boolean Matrix Logic Programming (BMLP) problem)
Let $\mathcal{P}$ be a datalog program containing a set of clauses with the predicate symbol $r$. The goal of Boolean Matrix Logic Programming (BMLP) is to find a boolean matrix R encoded by a datalog program such that $(\textbf{R})_{i,j}$ = 1 if $\mathcal{P}\models r(c_{i},c_{j})$ for constants $c_{i},c_{j}$ and $(\textbf{R})_{i,j}$ = 0 otherwise.
The $H_{2}^{2}$ program class when extended with functional symbols has Universal Turing Machine expressivity [31]. Consider a ground term $u$ that concatenates $n$ constant symbols $c_{1},c_{2},...,c_{n}$. We can represent $u$ using a $n$-bit binary code. When $c_{k}$ is in $u$ the $k$-th bit of the binary code is 1, otherwise, it is 0.
Figure 2: (a) Vector v encodes source chemical metabolites. A reaction is represented in two boolean matrices, $\textbf{R}_{1}$ and $\textbf{R}_{2}$. (b) We implemented a BMLP module BMLP-IE to compute $\textbf{v}^{*}$, the closure of pathway products, using boolean matrix operations addition, multiplication, equality and logic AND.
Two boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ map reactions to binary codes. The binary codes represent reactants and products. We implement a BMLP module called iterative extension (BMLP-IE) to perform the bottom-up evaluation of reachability in metabolic networks. BMLP-IE module (Fig 2) iteratively expands facts derivable from a partially grounded query $r(u,Y)$. Given that $\mathcal{P}$ contains $n$ constant symbols, we represent the constants in $u$ as a $1\times n$ row vector $\mathbf{v}$ (a $n$-bit binary code). In addition, a $1\times k$ boolean vector t is used to turn transitions in the OEN “on” or “off”. This allows us to temporarily modify the network connectivity without re-creating $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
5 Active learning
Our active learning system $BMLP_{active}$ selects experiments to minimise the expected value of a user-defined cost function and iteratively prunes hypotheses inconsistent with experimental outcomes. $BMLP_{active}$ uses the Minimal Description Length principle [7] and the compression score of a hypothesis $h$ [5] to optimise the posterior probability:
$compression(h,E)$ $=|E^{+}|-\frac{|E^{+}|}{pc_{h}}(size(h)+fp_{h})$
$p^{\prime}(h|E)$ $=\frac{2^{compression(h,E)}}{\sum_{h_{i}\in H}2^{compression(h_{i},E)}}$
$E$ are labelled examples and $E^{+}$ are positive examples. $pc_{h}$ and $fp_{h}$ are the number of positive predictions and false positive coverage. This compression function favours a general hypothesis with high coverage and penalises over-general hypotheses that incorrectly predict negative examples. $BMLP_{active}$ uses the following heuristic function from [5] to select an experiment from candidate experiments $T$ with approximated minimum expected cost:
$EC(H,T)\approx min_{t\in T}\,[C_{t}$ $+p(t)(mean_{t^{\prime}\in T-\{t\}}C_{t^{\prime}})J_{H_{t}}$
$+(1-p(t))(mean_{t^{\prime}\in T-\{t\}}C_{t^{\prime}})J_{\overline{H_{t}}}]$
$H_{t}$ and $\overline{H_{t}}$ are subsets of hypotheses $H$ consistent and inconsistent with $t$’s label. $J_{H_{t}}andJ_{\overline{H_{t}}}$ are calculated according to the entropy $-\sum_{h\in H}p^{\prime}(h|E)\,log_{2}(p^{\prime}(h|E))$ where $H$ is replaced by $H_{t}$ or $\overline{H_{t}}$. The probability $p^{\prime}(h|E)$ is calculated from the compression function. $p(t)$ is the probability that the label of the experiment $t$ is positive and is computed by the probability sum of consistent hypotheses $\sum_{h\in H_{t}}p^{\prime}(h|E)$. $C_{t}$ is a user-defined experiment cost function.
For some hypothesis space $H$ and background knowledge $BK$, $V_{s}$ is the version space consistent with $s$ training examples with the assumption that an active version space learner selects one instance per iteration. The shrinkage of the hypothesis space is $\frac{|V_{s+1}|}{|V_{s}|}$ after querying the label of the $s$-th instance. The minimal reduction ratio $p(x_{s},V_{s})$ is the minority ratio of the version space $V_{s}$ partitioned by an instance $x_{s}$.
$p(x_{s},V_{s})=\frac{min(\,|\{h\in V_{s}|\,h\cup BK\models x_{s}\}|,|\{h\in V_{s}|\,h\cup BK\not\models x_{s}\}|)}{|V_{s}|}$
The optional selection strategy is to select instances with minimal reduction ratios as close as possible to $\frac{1}{2}$ to eliminate up to $50\%$ of the hypothesis space [18]. A passive learner is considered a learner using random example selection since it does not have control over the training examples it uses.
Theorem 5.1
(Active learning sample complexity bound [1]) For some $\phi\in[0,\frac{1}{2}]$ and small hypothesis error $\epsilon>0$, if an active version space learner can select instances to label from an instance space $\mathcal{X}$ with minimal reduction ratios greater than or equal to $\phi$, the sample complexity $s_{active}$ of the active learner is
$s_{active}\leq\frac{\epsilon}{\epsilon+\phi}s_{passive}+c$ (1)
where c is a constant and $s_{passive}$ is the sample complexity of learning from randomly labelled instances.
Theorem 5.1 says that the number of instances needed to select by active learning should be some factor smaller than the number of randomly sampled examples given some predictive accuracy level. Recent work [1] has empirically supported Theorem 5.1. This means the biological experiment sequence needed to arrive at a finding would be shortened by active learning.
Labels of experiment instances are requested from a data source, e.g. a laboratory or an online dataset. We apply BMLP-IE to obtain predicted phenotypic effects for all combinations of candidate experiments and hypotheses which we are compared with experimental data. The compression of a hypothesis is calculated by its coverage for all unlabelled instances. Currently, we only consider labels from synthetic or online phenotype data. However, $BMLP_{active}$ can request data from a high-throughput experimental workflow to automate experiments.
6 Experiments
Figure 3: tyrB isoenzyme function recovery frequency. The experimental space had $(\binom{33}{2}+33)\times 7=3927$ instances for double gene-knockout synthetic data and single gene-knockout experimental data of key 33 genes in 7 conditions. The hypothesis space contained $(27\times 32+6\times 31)+2=1052$ candidate gene-enzyme associations for 27 single-function genes, 6 double-function genes, the tyrB original function and an empty hypothesis.
We focused on re-discovering the function associated with the key gene $tyrB$ in the Tryptophan biosynthesis pathways. Overlapping functions between two genes $tyrB$ and $aspC$ is a digenic interaction since they are responsible for producing an important amino acid. We removed the metabolic reaction associated with $tyrB$ from iML1515. This state-of-the-art GEM model iML1515 [19] was first represented as a datalog program and then as boolean matrices for computation. The random selection strategy randomly sampled $N$ instances from the instance space. $BMLP_{active}$ selected $N$ experiments from this instance space to actively learn from the hypothesis space. Both methods output the hypothesis with the highest compression.
In Fig 3 we observed that $BMLP_{active}$ provides higher information gain from each experiment and successfully guarantees recovery of the correct gene function with as few as 20 experiments. This demonstrates $BMLP_{active}$’s ability to discover digenic interactions in GEM. This result also shows that $BMLP_{active}$ significantly reduced the number of experiments needed compared to random experiment selection. Random sampling failed to eliminate competitive hypotheses beyond a certain point.
7 Conclusion and future work
We integrated abductive reasoning and Bayesian active learning in the new approach Boolean Matrix Logic Programming (BMLP), which utilises low-level boolean matrices encoded in Prolog to compute datalog programs. We applied our active learning system $BMLP_{active}$ to learn a key digenic function in a state-of-the-art genome-scale metabolic network. Though we focused on a reduced set of hypotheses in the experiment, the combinatorial space was sufficiently large that random experiment selection became non-viable as an experimentation strategy in a discovery process. The remarkable increase in efficiency with $BMLP_{active}$ with this task demonstrates it has potential even as the experimental design space grows exponentially.
Petri nets complement knowledge representation with dynamic analysis. However, the simulation of Petri nets in logic programming has been much less explored. Future work will explore the natural link between Petri nets and Probabilistic Logic Programming [9]. Transitions could have firing likelihood constraints and this uncertainty can be modelled by probabilistic logic programs.
{credits}
7.0.1 Acknowledgements
The first, third and fourth authors acknowledge support from the UKRI 21EBTA: EB-AI Consortium for Bioengineered Cells & Systems (AI-4-EB) award (BB/W013770/1). The second author acknowledges support from the UK’s EPSRC Human-Like Computing Network (EP/R022291/1), for which he acts as Principal Investigator.